# The Green Peace Dividend: the Effects of Militarization on Emissions and the Green Transition

### v1

@@ Abstract
This paper argues that military buildups lead to a significant rise in greenhouse gas emissions and can disrupt the green transition. Identifying military spending shocks, I use local projections to show that a percentage point rise in the military spending share leads to a 1-1.5% rise in total emissions, as well as a 1% rise in emission intensity. Using a dynamic production network model calibrated for the US, I find that a permanent shock of the same size would increase total emissions by between 0.36% and 1.81%, and emission intensity by between 0.22% and 1.5%. The model indicates that fossil fuel and energy-intensive firms experience a considerable expansion in response to such a shock, which could create political obstacles for the green transition. Similarly, investment in renewables and green R&D could be crowded out by defence spending, further hindering the energy transition. Policymakers can use carbon prices or green subsidies to counteract these effects, the latter likely being more efficient due to political and social constraints.
@@ 1 Introduction
After several decades of relative calm and peace, a flare-up of wars and a souring of international relations between the major powers of the world prompted economists and policymakers to redirect their attention toward understanding the effects of military activity.
@@
This paper aims to uncover, measure, and understand the effects of militarization on climate change mitigation. In particular, I investigate three main channels. First, military buildups can raise greenhouse gas emissions directly, through increased energy and fuel use and a larger volume of military procurements. Armed forces need more energy and fuel to run their military bases, planes, bombers, tanks, etc. at a higher intensity, which increases energy use, but they also need more of these weapons. Defence manufacturing is an energy-intensive part of the economy, but especially its inputs, like primary metals, rubber, or chemicals are very energy-intensive. Hence, increasing demand for these goods can drive up the emission intensity of an economy and consequently total emissions.
Second, military buildups can impede the green transition through indirect economic effects. The previously mentioned rise in demand for energy, fuel, and energy-intensive goods might lead to an expansion of fossil fuel and energy-intensive manufacturing industries. This could lead to a rise in the political power of these industries, which could complicate passing and implementing climate policies.
Third, such a demand shock can also lead to a crowd-out in green investment and innovation. Defence and fossil fuel industries will require large amounts of investment to expand capacity to the level where it can supply a substantially more active military. This can lead to a scarcity of investment funds, delaying and potentially disrupting the financing of the green transition.
I find direct evidence for the existence of the first mechanism. In my empirical analysis, I analyse country-year panel data, extract military spending shocks using the method proposed by Hamilton (2018), and estimate the dynamic effect of such a shock on emissions and emission intensity using panel local projections. The results indicate that total emissions rise by 1-1.5%, while emission intensity rises by about 1% in response to a percentage point rise in the military spending share.
@@
I find some indirect evidence for the second and third mechanism mentioned above. The model indicates that fossil fuel and energy-intensive industries experience a significant expansion, with investment in these sectors crowding out investment in other parts of the economy. Realignment of production towards dirty goods can, then, increase dirty lobbying efforts and reduce green lobbying. In addition, investment in the utilities sector experiences a slump in response to a military spending shock, which suggests that investment in new renewable energy sources will most likely also suffer. If R&D investment is similarly crowded out, then green innovation would also decline.
Policymakers can counteract these negative effects using either a "stick" or a "carrot" approach, or a combination of these. The former would entail a rise in carbon prices, reducing total GHG emissions. Unfortunately, it would most likely hit political and social constraints: military buildups and wars require welfare sacrifices in general, and carbon prices would most likely pile upon these pressures. The "carrot", investment and innovation subsidies, would likely be more popular, and could also be harmonised with defence goals such as energy security.
@@
This paper contributes to the literature on the links between military activity and environmental damages. It is closest to the sub-stream looking at the causal effects of military expenditures on greenhouse gas emissions. Bradford and Stoner (2017), Clark et al. (2010), Jorgenson and Clark (2015), and Bildirici (2016) conduct empirical analyses of the effect of military expenditures on emissions or emission intensity. They generally find positive effects, although Bradford and Stoner (2017) find a relatively small impact of the military expenditure share on emission intensity.
Most of these papers ignore the non-stationarity of the data (Bildirici (2016) tests and estimates the cointegrating relationship between military expenditures and emissions) and run simple panel regressions. Compared to them, I extract military spending shocks and calculate the dynamic effects of such a shock on emissions and emission intensity. In addition, I also use an economic, dynamic production network model to estimate these causal effects, as well as to uncover the underlying mechanisms. I also analyse the effect of such a shock on other long-run economic consequences, such as green investments and political economy consequences - I do not restrict my focus on GHG emissions.
@@
My paper also contributes to the literature on the economic effects of military buildups and wars. It is closest to two papers in particular. Neimark et al. (2024) and de Klerk et al. (2023) analyse the environmental costs of the Israel-Hamas war and the Ukraine war, respectively. Compared to them, I do not focus specifically on wars, and I also look at the longer-term effects of military buildups on the green transition.
Ramey (2011), Ramey and Shapiro (1998), Hall (1980), Barro (1981), Nakamura and Steinsson (2014), and Miyamoto et al. (2019) exploit military spending shocks to measure the effect of government spending on different economic outcomes - however, they do not analyse the link between military spending and emissions. Couttenier et al. (2022) and Korovkin and Makarin (2022) use production network models to estimate the economic effects of wars, in the context of the Maoist insurgency in India and the Ukraine war, respectively. I use methods similar to these papers to understand the effect of military activity, but I focus on climate-relevant impacts.
The paper proceeds in the following manner. I first discuss the empirical analysis in section 2, followed by the economic model in section 3. I sum up and interpret the results in section 4, where I discuss the Green Peace Dividend. In section 5, I discuss potential policy implications, after which I conclude.
@@ 2 Empirical analysis
In my empirical analysis, I aim to uncover how militarization affects greenhouse gas emissions, as well as the emission intensity of the economy. I proxy militarization with the military expenditure share of GDP.
@@ 2.1.1 Data
In my empirical analysis, I use annual GHG emissions data from EDGARv8.0. I am also using data from the Correlates of War (National Material Capabilities version 6.0) for energy consumption in tons of coal equivalent, as well as for military expenditures, military personnel, population, and urban population. National accounts data is from the Penn World Tables (version 10.01). I’m using data from V-Dem v13 measuring how close a given country is to the "ideal of electoral democracy" (Coppedge et al, 2023).
@@ 2.1.2 Method
To extract the causal effect of military spending on the outcomes of interest, I follow the approach in Bilal and Känzig (2024) and first extract military spending shocks, after which I estimate the dynamic effect of such a shock on the outcomes of interest using panel local projections (Jordà (2005), Jordà et al. (2020)). It is common in the literature to assume that military spending shocks occur due to geopolitical events, and not due to endogenous economic considerations. For example, Miyamoto et al. (2019) assume that "changes in military spending are exogenous to economic conditions". A similar argument can be found in Hall (1980).
@@
Where $y$ denotes either (log) emissions or (log) emission intensity, $\alpha$ is a country fixed effect, $\delta$ is a year fixed effect, $X_{i,t-1}$ is a vector of lagged controls for country i in year t, and $u$ is the regression error term.  [^4] The control variables include the first differences of logged GDP per capita, the export and import share of GDP, the urban population share, and a democracy variable (in levels), measuring the distance of a given country’s institutions in a given year from the ”ideal of electoral democracy” (Coppedge et al, 2023). I also add a second lag of GDP per capita. I also add $l$ lags of the first difference of the outcome variable to account for potential autocorrelation in the data - in practice, I set $l=2$. The IRF of the military spending shock is given by $IRF_{M}=\{\beta_{h}\}_{h=0}^{T}$, where I calculate the IRF up to horizon $T=15$.  [^5] The interpretation of $\beta_{h}$ is the following: if we increase the military spending share by 1 p.p., what’s the percentage change in the outcome relative to what it would have been without the shock, at horizon h? I use Driscoll and Kraay (1998) standard errors for inference, with two lags. In my IRFs, I show 68% and 95% confidence bands - as Ramey (2011) points out, 68% confidence intervals are common in the government spending multiplier literature.
@@ 2.2 Results
Note: the figure shows the IRF of the military spending share in response to a military spending shock corresponding to a doubling of the military spending share. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval, while the light grey area marks the 95% CI.
@@
Figure 4 shows the main results for the baseline, NATO sample. Both total emissions and emission intensity peak 6 years after the shock - emissions rise by 1.54%, while emission intensity rises by 1% in response to an increase in the military spending share of one percentage point. Two potential reasons why emission intensity increases, as discussed above, are, first, that the armed forces start using more fuel to run their weapons and bases, increasing the fuel use of the economy, and second, that the military industries are more emission-intensive than the rest of the economy, which means that a relative expansion of these industries raises the aggregate emission intensity of the economy.
@@
Total emissions rise both because emission intensity increases and because output rises. [^7] Note that, by definition: $Emissions=RealGDP\cdot EmissionIntensity$ So that: $d\ln(Emissions)=d\ln(RealGDP)+d\ln(EmissionIntensity)$ Figure 5 shows that output also rises in response to such a shock, by 1.11% after 4 years, which falls to approximately 0.67% after 6 years. This result is in line with the evidence in Ramey (2011) or Ilzetzki et al. (2013) concerning industrial countries (my main sample consists almost completely of industrialized countries). [^8] A CfM survey in July 2024 by Ilzetzki and Jain also found that most experts believed an increase in defence spending will boost economic growth in the EU over a 5-year horizon. https://www.cfmsurvey.org/survey-2024-07
@@
Emission intensity seems to increase somewhat significantly only in the horizon of 5-6 years. It is possible that the bulk of the military activity concentrates on this horizon after the shock - however, more evidence is needed to understand the dynamics of emission intensity following military buildups.
@@
Note that contemporary effects are virtually zero. A potential explanation for why the effect peaks much later than the year of the shock is that military buildups require time to materialize: weapon procurements can take years, and yet more time is needed for new weapons to be operationalized and new soldiers to be trained. [^9] This can, of course, happen in much shorter time periods, depending on the urgency of the military situation. During wartime, major industrial powers are capable of producing a very significant amount of matèriel in a very short time - an extreme example of this can be found in the Second World War and the performance of the formidable military industries of the United States, the Soviet Union and Nazi Germany. No such major events are included in my sample. Military buildups may require large-scale investments into military industries and their upstream suppliers, which can take time to materialize and become functional.
Also note, however, that the confidence bands are quite wide, so additional evidence would be important to confirm these results. The rise in total emissions is significant at the 5% significance level - however, the rise in emission intensity is not.
Figure 4: IRF - emissions and emission intensity
Note: the figure shows the IRFs of total emissions and emission intensity in response to a military spending shock corresponding to a doubling of the military spending share. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval, while the light grey area marks the 95% CI.
There is an effect which can reduce the measured impact on both emissions and emission intensity. Usually, part of the military budget is spent in foreign countries - both weapons (through arms trade) and energy could be bought externally. This means that part of the rise in emissions and emission intensity due to military buildups could show up in other countries’ data. This suggests that my results are conservative estimates of the true effect.
Figure 5: IRF - real GDP
Note: the figure shows the IRF of real GDP in response to a military spending shock corresponding to a doubling of the military spending share. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval, while the light grey area marks the 95% CI.
@@
Instead of emissions, one can also use energy use as an outcome variable to cross-check the results. Figure 6 shows that the results are very similar if we look at total energy use and energy intensity. The peak figures are slightly higher than in the case of emissions, but note that energy use does not necessarily translate one to one into emissions - the latter depends on the energy composition as well.
Figure 6: IRF - energy use and energy intensity
@@
Note: the figure shows the IRFs of total energy use and emission intensity in response to a military spending shock corresponding to a doubling of the military spending share. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval, while the light grey area marks the 95% CI.
@@
To check the generality of the previous results, I rerun the previous local projections with the extended sample of 38 countries, mentioned before.
Figure 7 shows the IRFs for this extended sample. The results are much weaker than those for the main sample: total emissions virtually do not change. This happens for two reasons: first, emission intensity does not change - in fact, it seems to fall in the long run (see second panel of figure 7). Second, output does not react to the shock, and only starts to increase after 10 years.
Figure 7: IRF - emissions and emission intensity (extended sample)
Note: the figure shows the IRFs of total emissions and emission intensity in response to a military spending shock corresponding to a doubling of the military spending share, in the extended sample. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval, while the light grey area marks the 95% CI.
Figure 8: IRF - real GDP (extended sample)
Note: the figure shows the IRF of real GDP in response to a military spending shock corresponding to a doubling of the military spending share, in the extended sample. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval, while the light grey area marks the 95% CI.
There might be several reasons behind these results. First, the domestic absorption of the additional countries in the sample might be lower than that of the NATO sample - countries such as Ireland likely do not source most of their military procurements domestically.
@@
Where $\mathbf{I}_{i,H}$ is equal to one if country i is classified as high-emission-intensity and zero if not, while $\mathbf{I}_{i,L}$ is equal to one if country i is classified as low-emission-intensity and zero if not. $\{\beta_{h,H}\}_{h=0}^{T}$ and $\{\beta_{h,L}\}_{h=0}^{T}$ are, then, the IRFs of the high- and low-emission-intensity countries. The high-emission-intensity IRFs can be seen on figure 9, while the low-emission-intensity IRFs are on figure 10.
Figure 9: IRF - emissions and emission intensity (high-emission-intensity countries)
@@
Note: the figure shows the IRFs of total emissions and emission intensity in response to a military spending shock corresponding to a doubling of the military spending share, in high-emission-intensity countries. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval, while the light grey area marks the 95% CI.
@@
Figure 10: IRF - emissions and emission intensity (low-emission-intensity countries)
@@
Note: the figure shows the IRFs of total emissions and emission intensity in response to a military spending shock corresponding to a doubling of the military spending share, in low-emission-intensity countries. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval, while the light grey area marks the 95% CI.
In high-emission-intensity countries, both emissions and emission intensity rise to a peak of 1% after 6 years, and the IRFs are similar to the main results concerning NATO countries in figure 4 (even though the samples are quite different). In low-emission-intensity countries, however, a military spending shock significantly reduces total emissions - by more than 3% after 3 years. It similarly reduces emission intensity, by almost 3%.
@@
Clearly, there is evidence of a differential response to a military spending shock in these two classes of countries. Hence, the emission and energy intensity of an economy may be important in determining whether such a shock actually leads to more GHG emissions or not. However, more research is needed into the exact reasons behind this result.

### v3

@@ Abstract
This paper argues that military buildups lead to a significant rise in greenhouse gas emissions and can disrupt the green transition. Identifying military spending shocks, I use local projections to show that a percentage point rise in the military spending share leads to a 0.9-2% rise in total emissions, as well as a 1% rise in emission intensity, depending on the economy’s overall emission intensity. Using a dynamic production network model calibrated for the US, I find that a permanent shock of the same size would increase total emissions by between 0.36% and 1.81%, and emission intensity by between 0.22% and 1.5%. The empirical analysis indicates that green patenting is reduced by 10-25% following such a shock, and the model suggests that investment in renewables could be crowded out by defence spending, hindering the energy transition. These effects can significantly raise climate damages, temperatures, and can lead to a quicker exhaustion of the carbon budget.
@@ 1 Introduction
After several decades of relative calm and peace in Europe, a flare-up of wars and a souring of international relations between the major powers of the world prompted economists and policymakers to redirect their attention toward understanding the effects of military activity.
@@
This paper aims to uncover, measure, and understand the effects of militarization on climate change mitigation. In particular, I investigate two main channels. First, military buildups can raise greenhouse gas emissions directly, through increased energy and fuel use and a larger volume of military procurements. Armed forces need more energy and fuel to run their military bases, planes, bombers, tanks, etc. at a higher intensity, which increases energy use, but they also need more of these weapons. Defence manufacturing is an energy-intensive part of the economy, but especially its inputs, such as primary metals, rubber, or chemicals are very energy-intensive. Hence, increasing demand for these goods can drive up the emission intensity of an economy and consequently total emissions.
Second, military buildups can impede the green transition through indirect economic effects. Such a demand shock can lead to a crowd-out in green investment and innovation. Defence and fossil fuel industries will require large amounts of investment to expand capacity to the level where it can supply a substantially more active military. This can lead to a scarcity of investment funds, delaying and potentially disrupting the financing of the green transition.
I find evidence for both of these mechanisms. In my empirical analysis, I use country-year panel data, extract military spending shocks using the method proposed by Hamilton (2018), and estimate the dynamic effect of such a shock on emissions and emission intensity using panel local projections. The results indicate that total emissions rise by 0.9-2%, while emission intensity rises by about 1% in response to a percentage point rise in the military spending share. [^2] Both the empirical analysis and the model results suggest that a rise in emission intensity, and not simply a rise in output due to a positive spending multiplier, is the main driving force behind these results. The result depends on how "dirty" is the given economy - that is, how emission-intensive the production structure is.
I also find evidence that military buildups reduce green innovation. The empirical analysis indicates that climate change mitigation and adaptation patents fall by 10-25% following a military expenditure shock, which can happen either due to a crowding-out effect or due to a general reduction in innovation.
@@
In addition, the model indicates that investment in the utilities sector might experience a slump in response to a temporary military spending shock, which suggests that investment in new renewable energy sources will most likely also suffer.
The result that the risk of military conflict can worsen climate change is important on its own because it shows that there is a "green peace dividend" - an additional cost of conflict. In section 5, I show that the measured effects can significantly raise climate damages, increase temperatures, and lead to the quicker exhaustion of the carbon budget. However, we also know that climate change can increase the likelihood of conflicts (see, e.g., Burke, Hsiang, and Miguel (2015) for a good summary of this literature). Hence, a positive feedback loop between conflict and climate change can increase the destructiveness of both.
Policymakers can counteract these negative effects using either a "stick" or a "carrot" approach, or a combination of these. The former would entail a rise in carbon prices, reducing total GHG emissions. Unfortunately, it would most likely hit political and social constraints: military buildups and wars require welfare sacrifices in general, and carbon prices would most likely pile upon these pressures. The "carrot", investment and innovation subsidies, would likely be more popular, and could be harmonised with defence goals such as energy security.
@@
This paper contributes to the literature on the links between military activity and environmental damages. It is closest to the sub-stream looking at the causal effects of military expenditures on greenhouse gas emissions. Bradford and Stoner (2017), Clark et al. (2010), Jorgenson and Clark (2015), Dong et al. (2024), and Bildirici (2016) conduct empirical analyses of the effect of military expenditures on emissions or emission intensity. They generally find positive effects, although Bradford and Stoner (2017) find a relatively small impact of the military expenditure share on emission intensity.
Most of these papers ignore the non-stationarity of the data (although Bildirici (2016) tests and estimates the cointegrating relationship between military expenditures and emissions) and run simple panel regressions. Compared to them, I extract military spending shocks and calculate the dynamic effects of such a shock on emissions and emission intensity. In addition, I also use a dynamic production network model to estimate these causal effects, as well as to uncover the underlying mechanisms. I also analyse the effect of such a shock on other long-run economic consequences, such as green investments or innovation - I do not restrict my focus on GHG emissions.
@@
My paper also contributes to the literature on the environmental effects of wars and conflict. It is closest to two papers in particular. Neimark et al. (2024) and de Klerk et al. (2023) analyse the environmental costs of the Israel-Hamas war and the Ukraine war, respectively. Compared to them, I do not focus specifically on wars, and I also look at the longer-term effects of military buildups on the green transition.
Ramey (2011), Ramey and Shapiro (1998), Auerbach and Gorodnichenko (2012), Hall (1980), Barro (1981), Nakamura and Steinsson (2014), and Miyamoto et al. (2019) exploit military spending shocks to measure the effect of government spending on different economic outcomes - however, they do not analyse the link between military spending and emissions. Couttenier et al. (2022) and Korovkin and Makarin (2022) use production network models to estimate the economic effects of wars, in the context of the Maoist insurgency in India and the Ukraine war, respectively. I use methods similar to these papers to understand the effect of military activity, but I focus on climate-relevant impacts.
The paper proceeds in the following manner. I first discuss the empirical analysis in section 2, followed by the economic model in section 3. I sum up and interpret the results in section 4, where I discuss the Green Peace Dividend and potential policy implications, after which I conclude.
@@ 2 Empirical analysis
In my empirical analysis, I aim to uncover how militarization affects greenhouse gas emissions, as well as green innovation. I proxy militarization with the military expenditure share of GDP.
@@ 2.1.1 Data
In my empirical analysis, I use annual GHG emissions data from EDGARv8.0. I also use data from the Correlates of War (National Material Capabilities version 6.0) for energy consumption in tons of coal equivalent, as well as for military expenditures, military personnel, crude steel production, population, and urban population. National accounts data is from the Penn World Tables (version 10.01). I’m using data from V-Dem v13 measuring how close a given country is to the "ideal of electoral democracy" (Coppedge et al, 2023). I use OECD data for total and environmental patents.
@@ 2.1.2 Method
To extract the causal effect of military spending on the outcomes of interest, I follow the approach in Bilal and Känzig (2024) and first extract military spending shocks using a Hamilton (2018) filter, after which I estimate the dynamic effect of such a shock on the outcomes of interest using panel local projections (Jordà (2005), Jordà et al. (2020)). I assume that these shocks are exogenous, as sudden changes in defence spending are most likely to follow strategic, and not economic, considerations. It is common in the literature to assume that military spending shocks occur due to geopolitical events, and not due to endogenous economic considerations. For example, Miyamoto et al. (2019) assume that "changes in military spending are exogenous to economic conditions". A similar argument can be found in Hall (1980).
@@
Where $y$ denotes either (log) emissions or (log) emission intensity, $\alpha$ is a country fixed effect, $\delta$ is a year fixed effect, $X_{i,t-1}$ is a vector of lagged controls for country i in year t, and $u$ is the regression error term.  [^5] The control variables include the first differences of logged GDP per capita, exports and imports, and a democracy variable (in levels), measuring the distance of a given country’s institutions in a given year from the ”ideal of electoral democracy” (Coppedge et al, 2023). I also add a second lag of GDP per capita. I also add $l$ lags of the first difference of the outcome variable to account for potential autocorrelation in the data - in practice, I set $l=2$. The IRF of the military spending shock is given by $IRF_{M}=\{\beta_{h}\}_{h=0}^{T}$, where I calculate the IRF up to horizon $T=15$.  [^6] The interpretation of $\beta_{h}$ is the following: if we increase the military spending share by 1 p.p., what’s the percentage change in the outcome relative to what it would have been without the shock, at horizon h? I use Driscoll and Kraay (1998) standard errors for inference, with two lags. In my IRFs, I show 68% confidence bands - as Ramey (2011) points out, 68% confidence intervals are common in the government spending multiplier literature.
@@ 2.2 Results
First, I will be looking at a military spending share shock’s effect on emissions, after which I will look at its effects on green patenting.
Emissions
@@
Note: the figure shows the IRF of the military spending share in response to a military spending shock corresponding to a 1 p.p. rise in the military spending share. The vertical axis shows percent changes. The dark grey area marks the 68% confidence interval.
@@
Figure 4 shows the main results for the baseline, NATO sample. Both total emissions and emission intensity (panels (a) and (b)) peak approximately 6 years after the shock - emissions rise by 1.94%, while emission intensity rises by 0.9% in response to an increase in the military spending share of one percentage point. Two potential reasons why emission intensity increases, as discussed above, are, first, that the armed forces start using more fuel to run their weapons and bases, increasing the fuel use of the economy, and second, that the military industries are more emission-intensive than the rest of the economy, which means that a relative expansion of these industries raises the aggregate emission intensity of the economy.
@@
Total emissions rise both because emission intensity increases and because output rises. [^8] Note that, by definition: $Emissions=RealGDP\cdot EmissionIntensity$ So that: $d\ln(Emissions)=d\ln(RealGDP)+d\ln(EmissionIntensity)$ Figure 5 shows that output also rises in response to such a shock, by 1.11% after 4 years, which falls to approximately 0.67% after 6 years. This result is in line with the evidence in Ramey (2011), Auerbach and Gorodnichenko (2012), and Ilzetzki et al. (2013) concerning industrial countries (my main sample consists almost completely of industrialized countries). [^9] A CfM survey in July 2024 by Ilzetzki and Jain also found that most experts believed an increase in defence spending will boost economic growth in the EU over a 5-year horizon. https://www.cfmsurvey.org/survey-2024-07
@@
Emission intensity seems to increase significantly only in the horizon of 6-7 years. The bulk of the military activity may concentrate on this horizon after the shock - however, more evidence is needed to understand the dynamics of emission intensity following military buildups. Nevertheless, the rise in emission intensity makes up almost half of the peak increase in total emissions, suggesting that the rise in emissions is not simply due to a positive government spending multiplier driving up output.
I will discuss the mechanism behind the effect of military spending on emissions in section 3. However, figure 6 shows that a defence spending shock increases crude steel production in the baseline sample countries by more than 2% after 6 years, which indicates that heavy manufacturing experiences a significant boost (even relative to real GDP, which rises by approximately 1% at this horizon). This provides evidence that military spending shocks target energy-intensive industries, such as steel manufacturing.
@@
Note that contemporary effects are virtually zero. A potential explanation for why the effect peaks much later than the year of the shock is that military buildups require time to materialize: weapon procurements can take years, and yet more time is needed for new weapons to be operationalized and new soldiers to be trained. [^10] This can, of course, happen in much shorter time periods, depending on the urgency of the military situation. During wartime, major industrial powers are capable of producing a very significant amount of matèriel in a very short time - an extreme example of this can be found in the case of the Second World War and the performance of the formidable military industries of the United States, the Soviet Union and Nazi Germany. See, for example, Ilzetzki (2024). No such major events are included in my sample. Military buildups may require large-scale investments into military industries and their upstream suppliers, which can take time to materialise and become functional.
@@
Instead of emissions, one can also use energy use as an outcome variable to cross-check the results - see panels (c) and (d) on figure 4. The results are very similar if we look at total energy use and energy intensity. Note that energy use does not necessarily translate one-to-one into emissions - the latter depends on the energy composition as well. Using data on the energy mix [^11] Energy Institute - Statistical Review of World Energy (2024) – with major processing by Our World in Data. “Biofuels consumption” [dataset]. Energy Institute, “Statistical Review of World Energy” [original data]., I find that the share of fossil fuels out of total energy consumption (in TWh) increases by 0.3 percentage points at the peak, driven mainly by higher natural gas consumption, suggesting that energy use becomes slightly "dirtier".
Figure 4: Baseline results - emissions and energy
@@
Note: the figure shows the IRFs of total emissions, energy use, emission intensity, and energy intensity in response to a military spending shock corresponding to a 1 p.p. rise in the military spending share. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
Note that there is an effect which can reduce the measured impact on both emissions and emission intensity. Usually, part of the military budget is spent in foreign countries - both weapons (through arms trade) and energy could be bought externally. Indeed, a large share of EU defence procurement spending goes to non-EU suppliers - between June 2022 and June 2023, 78% of EU defence procurement spending went to non-EU suppliers (see the Draghi report (2024)). This means that part of the rise in emissions and emission intensity due to military buildups could show up in other countries’ data, which suggests that my results are conservative estimates of the true effect.
Figure 5: Real GDP
Note: the figure shows the IRF of real GDP in response to a military spending shock corresponding to a 1 p.p. rise in the military spending share. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
Figure 6: Crude steel production
Note: the figure shows the IRF of crude steel production in response to a military spending shock corresponding to a 1 p.p. rise in the military spending share. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
@@
To check the generality of the previous results, I rerun the previous local projections with the extended sample of 38 countries, mentioned before. The results are much weaker than those for the main sample: total emissions virtually do not change. There might be several reasons behind this. First, the domestic absorption of the additional countries in the sample might be lower than that of the NATO sample - countries such as Ireland likely do not source most of their military procurements domestically.
@@
Where $\mathbf{I}_{i,H}$ is equal to one if country i is classified as high emission intensity and zero if not, while $\mathbf{I}_{i,L}$ is equal to one if country i is classified as low emission intensity and zero if not. $\{\beta_{h,H}\}_{h=0}^{T}$ and $\{\beta_{h,L}\}_{h=0}^{T}$ are, then, the IRFs of the high and low emission intensity countries. The high-emission-intensity IRFs can be seen on figure 7, while the low-emission-intensity IRFs are on figure 9.
Figure 7: High emission intensity countries - results
@@
Note: the figure shows the IRFs of total emissions and emission intensity in response to a military spending shock corresponding to a 1 p.p. rise in the military spending share, in high emission intensity countries. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
@@
Figure 8: High emission intensity countries - output and steel production
@@
Note: the figure shows the IRFs of real GDP and crude steel production in response to a military spending shock corresponding to a 1 p.p. rise in the military spending share, in high emission intensity countries. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
In high-emission-intensity countries, emissions rise by approximately 0.9% 6 years after the shock, while emission intensity rises to a peak of more than 1%. This indicates that, in this sample, military spending shocks drive emissions up solely by making the economy "dirtier": figure 8 shows that real GDP does not change significantly following a shock and that crude steel production rises by 1% at a horizon of 5 years.
In low-emission-intensity countries, however, a military spending shock significantly reduces total emissions - by 3% after 3 years (see figure 9). It similarly reduces emission intensity, by 2%. Figure 10 suggests that both output and crude steel production fall in this sample, which could suggest that, instead of the production structure being "cleaner" in this sample, the results could be driven by low domestic absorption (redirecting demand from domestically produced goods to foreign-produced weapons).
Figure 9: Low emission intensity countries - results
Note: the figure shows the IRFs of total emissions and emission intensity in response to a military spending shock corresponding to a 1 p.p. rise in the military spending share, in low emission intensity countries. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
Figure 10: Low emission intensity countries - output and steel production
Note: the figure shows the IRFs of real GDP and crude steel production in response to a military spending shock corresponding to a 1 p.p. rise in the military spending share, in low emission intensity countries. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
@@
Clearly, there is evidence of a differential response to a military spending shock in these two classes of countries. Hence, the emission and energy intensity of an economy may be important in determining whether such a shock actually leads to more GHG emissions or not. However, more research is needed into the exact reasons behind this result - for example, differences may be due to higher domestic absorption of defence expenditures in high emission intensity countries.
Innovation
In order to understand the effect of a military spending shock on green innovation, I look at the effects on patents linked to climate change mitigation and adaptation using the local projection setup (2). However, several countries in my sample produce few or no patents targeting these areas - hence, I focus on countries in the 38-country sample that are the most innovative in this area. That is, I calculate the time series averages of green patents in each country between 1978 and 2016 and only use the countries which are above the median of these averages. [^12] These countries are: Australia, Austria, Belgium, Canada, Chile, Denmark, Finland, France, Germany, Italy, Japan, Korea, the Netherlands, Sweden, Switzerland, Taiwan, the UK and the US. Then, I run (2) for these countries, for the period between 1978-2016.
The results can be seen on figure 11. A military spending share shock reduces environmental patents by 10% contemporaneously and after a year, following which it returns to the trend (panel (a)). This is not due to a general negative effect on patenting: panel (b) shows that total patents increase significantly, by more than 20% following a defence shock.
Total patents could increase due to two effects: first, because military R&D boosts innovation directly, and second, because defence R&D could "crowd-in" private innovation (see Moretti, Steinwender, and Van Reenen (2021) on the "crowd-in" effects of defence R&D). However, the results suggest that as overall innovation receives a boost, green innovation is crowded out - potentially due to the lack of strong links between military and green research.
Figure 11: Patenting - green and total
Note: the figure shows the IRFs of total patents and green patents in response to a military spending shock corresponding to a 1 p.p. rise in the military spending share. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
I further restrict the sample to look at high-patenting NATO countries - results can be seen on figure 12. Both environmental patenting, as well as total patents experience a sharp drop following a 1 p.p. rise in the military expenditure share. Green patenting falls relatively more than total patents (by 25% vs. 17%), suggesting that green innovation suffers disproportionately in NATO countries following a defence spending shock. Nevertheless, the results imply that military expenditures might have a general negative effect on innovation and not simply crowd out green patenting in this smaller sample. [^13] Note that Moretti et al. (2021) find that defence-related R&D expenditures increase industry-level private R&D expenditures and, hence, total industry-level R&D expenditures. Hence, a possible reason why total patenting falls in my smaller, NATO sample is that the identified military expenditure share shocks might contain small R&D shares.
Figure 12: Patenting - green and total (NATO countries)
Note: the figure shows the IRFs of total patents and green patents in response to a military spending shock corresponding to a 1 p.p. rise in the military spending share, in NATO countries. The outcome is in terms of percentage changes. The dark grey area marks the 68% confidence interval.
As the results - especially for the larger sample of innovative countries - are quite noisy with wide confidence bands, further research is necessary to understand the effect of defence expenditure shocks on green innovation.